# PostEdit: Posterior Sampling for Efficient Zero-Shot Image Editing

### v1

@@ Abstract
In the field of image editing, three core challenges persist: controllability, background preservation, and efficiency. Inversion-based methods rely on time-consuming optimization to preserve the features of the initial images, which results in low efficiency due to the requirement for extensive network inference. Conversely, inversion-free methods lack theoretical support for background similarity, as they circumvent the issue of maintaining initial features to achieve efficiency. As a consequence, none of these methods can achieve both high efficiency and background consistency. To tackle the challenges and the aforementioned disadvantages, we introduce PostEdit, a method that incorporates a posterior scheme to govern the diffusion sampling process. Specifically, a corresponding measurement term related to both the initial features and Langevin dynamics is introduced to optimize the estimated image generated by the given target prompt. Extensive experimental results indicate that the proposed PostEdit achieves state-of-the-art editing performance while accurately preserving unedited regions. Furthermore, the method is both inversion- and training-free, necessitating approximately 1.5 seconds and 18 GB of GPU memory to generate high-quality results. Code is available at https://github.com/TFNTF/PostEdit.
@@
Inversion-based approaches Song et al. (2021a); Mokady et al. (2023); Wu & la Torre (2022); Huberman-Spiegelglas et al. (2024) progressively add noise to a clean image and then remove the noise conditioned on a given target prompt, utilizing large text-to-image diffusion models (i.e. Stable Diffusion Rombach et al. (2022)), to obtain the edited image. However, directly inverting the diffusion sampling process (e.g., DDIM Song et al. (2021a)) for reconstruction introduces bias from the initial image due to errors accumulated by an unconditional score term, as discussed in classifier-free guidance (CFG) Ho & Salimans (2022) and proven in App. A.1. Consequently, the editing quality of inversion-based methods is primarily constrained by the similarity in unedited regions. Several approaches address this issue by optimizing the text embedding Wu et al. (2023), employing iterative guidance Kim et al. (2022); Garibi et al. (2024), or directly modifying attention layers Hertz et al. (2023); Mokady et al. (2023); Parmar et al. (2023) to mitigate the bias introduced by the unconditional term. However, the necessity of adding and subsequently removing noise predicted by a network remains unavoidable, thereby significantly constraining their efficiency. Recent methods Starodubcev et al. (2024); Li & He (2024); Kim et al. (2024) attempt to enhance the accuracy of the iterative sampling process by training an invertible consistency trajectory, following the distillation process in the consistency models (CM) Song et al. (2023); Salimans & Ho (2022); Song & Dhariwal (2024); Luo et al. (2023b). Although this approach significantly reduces the accumulation errors from the unconditional term, it cannot eliminate them. Moreover, the editing performance is sensitive to the hyperparameters (i.e., the fixed boundary timesteps of multi-step consistency models), and the training process generally demands hundreds of GPU hours.
@@
Another category of methods Brooks et al. (2023); Mou et al. (2024); Ye et al. (2023); Guo et al. (2024); Li et al. (2023); Wang et al. (2024) is inversion-free, which omits the inversion process and consequently significantly decreases the inference time. Specifically, these methods follow the approach of ControlNet Zhang et al. (2023b) to train a network that embeds additional conditional features, such as hand-drawn outlines, depth maps, and canny maps. However, the works based on ControlNet like T2I-Adapter Mou et al. (2024) highly rely on the accuracy of the input guidance structure while most applications relating to ControlNet are customization. Some other works Zhang et al. (2023a), Zhang et al. (2024b), Hui et al. (2024) employ a diffusion model trained on synthetic edited images, producing edited images in a supervised manner. This methodology obviates the need for inversion process during the sampling stage. Moreover, there is a training-free method to satisfy inversion-free requirement Xu et al. (2024). It adopts specific settings of the DDIM solver to leverage the advantages of CM to ensure the editing quality. Although these recent works can achieve fast sampling and accurate editing, the aforementioned problem remains unsolved since the diffusion sampling process Ho et al. (2020); Song et al. (2021a; b) is necessary. Therefore, all the inversion-free methods cannot circumvent the accumulation errors caused by the unconditional score term in CFG.
Figure 1: Comparison of Different Methods. The process of inversion-based method is shown on the top-left part. It progressively add noise to a clean image by applying the noise estimated by the network and then obtains the edited image by a denoise process. There is an optimization process for maintaining the background similarity. The too-right part is the methods with an inversion-free process, which adds noise randomly and operates training process on network to get satisfied output. Our method is designed to become inversion-free and training-free as shown in the bottom.
@@
In this work, we present an inversion- and training-free method called PostEdit to optimize the accumulated errors of the unconditional term in CFG based on the theory of posterior sampling Kawar et al. (2022),Chung et al. (2023), Zhang et al. (2024a), Lugmayr et al. (2022), Zhu et al. (2023), Song et al. (2021b). To reconstruct and edit an image $\boldsymbol{x}_{0}$, we make use of a measurement term $\boldsymbol{y}$ which contains the features of the initial image, and supervise the editing process by the posterior log-likelihood density $\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{x}_{t}|\boldsymbol{y})$. With this term, we can estimate the target image through progressively sampling from the posterior $p(\boldsymbol{x}_{t}|\boldsymbol{y})$ referring to the Bayes rule. The above process is reasonable since the inverse problems of probabilistic generative models are ubiquitous in generating tasks, which are trained to learn scores to match gradients of noised data distribution (log density), and this process is also called score matching Song & Ermon (2020), Song & Ermon (2019), Karras et al. (2022) and Karras et al. (2024). $\boldsymbol{y}$ is defined according to the following inverse problem
@@ 2.1 Score-Based Diffusion Models
A continuous diffusion process as shown in Song et al. (2021b) is adopted in this paper to sample the estimated initial image $\boldsymbol{\hat{x}}_{0}$. Specifically, the forward diffusion process can be modeled as the solution to an Itô SDE:
@@
The ODE is able to compute the exact likelihood of any input data by leveraging the connection to neural ODEs Chen et al. (2018). In order to approximate the log density of noised data distribution $\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$ for each sampling step, a network Song et al. (2021b) $\boldsymbol{s}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t)$ is trained to learn the corresponding log density
@@ 2.2 DDIM Solver And Consistency Models
$\boldsymbol{\hat{x}}_{0}=\frac{\boldsymbol{x}_{t}-\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t)}{\sqrt{\alpha_{t}}}$ (7)
@@ 2.3 Posterior Sampling in Diffusion Models
$\boldsymbol{x}_{t-1}=f\left(\boldsymbol{x}_{t},\boldsymbol{\hat{x}}_{0},\boldsymbol{\epsilon}\right)+\eta\nabla_{\boldsymbol{x}_{t}}\left\|y-\mathcal{A}(\boldsymbol{\hat{x}}_{0})\right\|^{2}_{2},\ \ \boldsymbol{\epsilon}\sim\mathcal{N}\left(0,\boldsymbol{I}\right),$ (12)
@@ 3 Method
We propose a new sampling process equipped with a designed optimization process to improve the quality of image reconstruction and editing, and the adopted SDE/ODE solver is based on DDIM shown in Eq. 6. Denote $\boldsymbol{z}\sim\mathcal{E}(\boldsymbol{x}_{0})$, $\boldsymbol{z}\in\mathbb{R}^{p}$ where $\mathcal{E}$ is an encoder and $\boldsymbol{x}_{0}$ is an initial image. Our method operates in latent space and leverages the theory of the posterior sampling to correct the bias from the initial features and introduce the target concepts. The core insight is using the measurements $\boldsymbol{y}$, estimated $\boldsymbol{z}_{0}$ and Langevin dynamics as the optimization terms to correct the errors of the sampling process. The importance of reconstruction and the algorithm are introduced specifically in (Sec. 3.1). The implementation details of the editing process are illustrated in detail (Sec. 3.2). Our method takes around 1.5 seconds and 18 GB memory costs on a single NVIDIA A100 GPU.
@@ 3.1 Posterior Sampling for Image Reconstruction
The quality of reconstruction is a crucial indicator for evaluating the editing capabilities of a method. To preserve the features of the background (areas unaffected by the target prompt), Mokady et al. (2023) introduces a technique for fine-tuning the text embedding to mitigate errors caused by the null text term, as demonstrated in App. A.1. However, this approach is time-consuming, and there is a pressing need to enhance editing performance. To address this challenge, we propose a method that enables a fast and accurate reconstruction and editing process.
@@
Specifically, there are four steps in our method: (1) We add noise to $\boldsymbol{z}_{0}$ by DDPM noise schedule till $\boldsymbol{z}_{T}\sim\mathcal{N}(0,\boldsymbol{I})$. Compared to the DDIM inversion process applied in Mokady et al. (2023), the added noise is sampled from $\mathcal{N}(0,\boldsymbol{I})$ so that it is not required to inference with networks many times. Consequently, this process is inversion-free and significantly decreases the time costs. (2) Existing SDE/ODE solvers, such as DDIM and LCM, are applied to obtain an estimated $\boldsymbol{\hat{z}}_{0}$. (3) In order to make $\boldsymbol{\hat{z}}_{0}$ become more consistent with the features of background and target prompt of the original image, it is optimized by two $\mathcal{L}_{2}$ norm terms relating to the defined measurement $\boldsymbol{y}$ and $\boldsymbol{\hat{z}}_{0}$ respectively. Further, the Langevin dynamics is applied to circumvent solutions of local optimal. (4) Finally, by progressively operating the above process to update the mean and variance of the Gaussian distribution in step (1) according to a time schedule, we can obtain the $\boldsymbol{z}_{0}^{*}$ with consistent initial features and accurate target characteristics respectively when $T$ is converged to 00. The presented algorithm corresponding to the above process is shown in detail in Fig. 2 for image reconstruction or editing and Alg. A.2 for image reconstruction.
@@
Our method requires a large text-to-image diffusion model as input and we select Stable Diffusion (SD) Rombach et al. (2022). Given that SD is trained on the dataset containing billions of images, the generated result has strong randomness relating to the same prompt. Moreover, reconstructing an image with a given text prompt starting from $\mathcal{N}\left(0,\boldsymbol{I}\right)$ usually gets poor results for SD due to the bias caused by the unconditional term in CFG. Therefore, if we directly apply posterior sampling strategy shown in Eq. 12 to acquire $\boldsymbol{z}_{0}$ from $\boldsymbol{z}_{T}\sim\mathcal{N}(0,\boldsymbol{I})$ by leveraging SD to inference noise, $\boldsymbol{\hat{z}}_{0}$ differs greatly from the ground truth $\boldsymbol{z}_{0}$. Conversely, this sampling process has good reconstruction performance when it leverages the diffusion model trained on small datasets, for example, FFHQ Karras et al. (2019) and ImageNet Deng et al. (2009) as shown in Chung et al. (2023); Zhang et al. (2024a). We experimentally discover that the gap between these two kinds of models is the inconsistent layouts of each estimated image $\boldsymbol{\hat{z}}_{0}$, while the features of the target prompts are successfully introduced into the generated $\boldsymbol{\hat{z}}_{0}$. Specifically, The layouts of $\boldsymbol{\hat{z}}_{0}$ generated by the scores inferred by the networks trained on FFHQ and SD for intermediate timesteps are shown in App. A.5. Therefore, due to the editing and reconstruction trade-off issue, it is much more challenging for high-quality image editing and reconstruction by leveraging large text-to-image models.
@@
to calculate the scores towards to $\boldsymbol{z}_{0}$ straightly inspired by Chung et al. (2023) and Zhang et al. (2024a). The measurement settings for image reconstruction are listed in the App. A.4
@@
$\boldsymbol{z}_{t-1}\sim\mathbb{E}_{\boldsymbol{z}_{0}^{w}}{\mathcal{N}}\left(\boldsymbol{z}_{0}^{w},\sigma_{t-1}^{2}\boldsymbol{I}\right)$ (15)
@@
satisfies the time marginal distribution conditioned on $p\left(\boldsymbol{z}_{t-1}|\boldsymbol{y}\right)$, where $\boldsymbol{z}_{0}^{w}$ is obtained from Eq. 13. (Proof is shown in Appendix. A.3)
@@
Propsition 2 ensures that $\boldsymbol{z}_{t-1}$ sampled from the Gaussian distribution with mean of $\boldsymbol{z}_{0}^{w}$ and variance of $\sigma_{t-1}^{2}$ respectively still satisfies the constraint of the posterior sampling Eq. 10. Therefore, we can present the following scheme to optimize the estimated $\boldsymbol{\hat{z}}_{0}$ and run Langevin dynamics Welling & Teh (2011)
@@
$\boldsymbol{z}_{0}^{(k+1)}=(1-w)\cdot\boldsymbol{z}_{0}^{(k)}+w\cdot\boldsymbol{z}_{in}-h\cdot\left(\frac{\|z_{0}^{(k)}-z_{0}\|^{2}}{2\sigma_{t}^{2}}+\frac{\|\mathcal{A}\left(\boldsymbol{z}_{0}^{(k)}\right)-\boldsymbol{y}\|^{2}}{2m^{2}}\right)+\sqrt{2h}\epsilon,\ \ \epsilon\sim\mathcal{N}(\boldsymbol{0},\mathcal{I}).$ (16)
@@
Experimental results in Fig 10 in Appendix. A.7 reflect that we can add noise randomly or completely sample from $\mathcal{N}\left(0,\boldsymbol{I}\right)$ to obtain a high quality reconstruction outcomes without any tuning process by our method. Eq. 16 is reasonable since the two terms that multiplied by the step size $h$ have the same descent direction towards to the ground truth $\boldsymbol{z}_{0}$. Additionally, the Langevin dynamics is applied to search for solutions with global optimal. Since the step (1) shown in Fig. 2 is different from the process of transitioning from the initial image to noise , which involves adding noise $\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{z}_{t},t,c_{ini}\right)$ inferred by the network at each step Mokady et al. (2023) (where $c_{ini}$ represents the prompt describing the content of the initial image). PostEdit is much more efficient as mentioned before. Considering the fact that there is no information relating to the initial image added to the noised distribution, the term containing the measurement $\boldsymbol{y}$ defined in Eq. 16 is introduced to correct the errors from the initial features caused by the unconditional term shown in CFG, ensuring background similarity.
The detailed process is outlined in App. A.2. As mentioned earlier, Proposition 1 is introduced due to the flexible layouts of the images generated by stable diffusion, which are not essential for the editing task. Additionally, the implementation details are provided in Sec. 3.2, following the introduction of more parameters.

### v2

@@ Abstract
In the field of image editing, three core challenges persist: controllability, background preservation, and efficiency. Inversion-based methods rely on time-consuming optimization to preserve the features of the initial images, which results in low efficiency due to the requirement for extensive network inference. Conversely, inversion-free methods lack theoretical support for background similarity, as they circumvent the issue of maintaining initial features to achieve efficiency. As a consequence, none of these methods can achieve both high efficiency and background consistency. To tackle the challenges and the aforementioned disadvantages, we introduce PostEdit, a method that incorporates a posterior scheme to govern the diffusion sampling process. Specifically, a corresponding measurement term related to both the initial features and Langevin dynamics is introduced to optimize the estimated image generated by the given target prompt. Extensive experimental results indicate that the proposed PostEdit achieves state-of-the-art editing performance while accurately preserving unedited regions. Furthermore, the method is both inversion- and training-free, necessitating approximately 1.5 seconds and 18 GB of GPU memory to generate high-quality results.
@@
Inversion-based approaches Song et al. (2021a); Mokady et al. (2023); Wu & la Torre (2022); Huberman-Spiegelglas et al. (2024) progressively add noise to a clean image and then remove the noise conditioned on a given target prompt, utilizing large text-to-image diffusion models (i.e. Stable Diffusion Rombach et al. (2022)), to obtain the edited image. However, directly inverting the diffusion sampling process (e.g., DDIM Song et al. (2021a)) for reconstruction introduces bias from the initial image due to errors accumulated by an unconditional score term, as discussed in classifier-free guidance (CFG) Ho & Salimans (2022) and proven in App. A.14. Consequently, the editing quality of inversion-based methods is primarily constrained by the similarity in unedited regions. Several approaches address this issue by optimizing the text embedding Wu et al. (2023), employing iterative guidance Kim et al. (2022); Garibi et al. (2024), or directly modifying attention layers Hertz et al. (2023); Mokady et al. (2023); Parmar et al. (2023) to mitigate the bias introduced by the unconditional term. However, the necessity of adding and subsequently removing noise predicted by a network remains unavoidable, thereby significantly constraining their efficiency. Recent methods Starodubcev et al. (2024); Li & He (2024); Kim et al. (2024) attempt to enhance the accuracy of the iterative sampling process by training an invertible consistency trajectory, following the distillation process in the consistency models (CM) Song et al. (2023); Salimans & Ho (2022); Song & Dhariwal (2024); Luo et al. (2023b). Although this approach significantly reduces the accumulation errors from the unconditional term, it cannot eliminate them. Moreover, the editing performance is sensitive to the hyperparameters (i.e., the fixed boundary timesteps of multi-step consistency models), and the training process generally demands hundreds of GPU hours.
@@
Another category of methods Brooks et al. (2023); Mou et al. (2024); Ye et al. (2023); Guo et al. (2024); Li et al. (2023); Wang et al. (2024) is inversion-free and thus significantly decreases the inference time. The general idea is to train networks to learn to embed the given conditions into the noisy-to-image diffusion process. For example, ControlNet Zhang et al. (2023b) and T2I-Adapter Mou et al. (2024) train an extra network to encode the image-shaped conditions, e.g., depth maps, canny maps. However, these works highly rely on the accuracy of the input guidance structure, while most applications related to ControlNet involve customization. Some other works Zhang et al. (2023a), Zhang et al. (2024b), Hui et al. (2024) employ a diffusion model trained on synthetic edited images, producing edited images in a supervised manner. This methodology obviates the need for inversion process during the sampling stage. Moreover, there is a training-free method to satisfy inversion-free requirement Xu et al. (2024). It adopts specific settings of the DDIM solver to leverage the advantages of CM to ensure the editing quality. Although these recent works can achieve fast sampling and accurate editing, the aforementioned problem remains unsolved since the diffusion sampling process Ho et al. (2020); Song et al. (2021a; b) is necessary. Therefore, all the inversion-free methods cannot circumvent the accumulation errors caused by the unconditional score term in CFG.
Figure 1: Different Image Editing Schemes. The inversion-based method, illustrated in the top-left section, involves adding noise from a pre-trained network to a clean image. It then denoises the image based on a target prompt, though it requires time-consuming tuning to ensure background preservation. The top-right section discusses training-based, inversion-free methods, which train a learnable model to achieve satisfactory results but have limited generalization capabilities. Our approach, outlined in the bottom section, is both inversion-free and training-free.
@@
In this work, we present an inversion- and training-free method named PostEdit to optimize the accumulated errors of the unconditional term in CFG based on the theory of posterior sampling Kawar et al. (2021; 2022); Chung et al. (2023); Zhang et al. (2024a; a); Lugmayr et al. (2022); Zhu et al. (2023); Song et al. (2021b). To reconstruct and edit an image $\boldsymbol{x}_{0}$, we adopt a measurement term $\boldsymbol{y}$ which contains the features of the initial image, and supervise the editing process by the posterior log-likelihood density $\nabla_{\boldsymbol{x}_{t}}\log p(\boldsymbol{x}_{t}|\boldsymbol{y})$. With this term, we can estimate the target image through progressively sampling from the posterior $p(\boldsymbol{x}_{t}|\boldsymbol{y})$ referring to the Bayes rule. The above process is reasonable since the inverse problems of probabilistic generative models are ubiquitous in generating tasks, which are trained to learn scores to match gradients of noised data distribution (log density), and this process is also called score matching Song & Ermon (2020), Song & Ermon (2019), Karras et al. (2022) and Karras et al. (2024). $\boldsymbol{y}$ is defined according to the following inverse problem
@@ 2.1 Score-Based Diffusion Models
We follow the continuous diffusion trajectory  Song et al. (2021b) to sample the estimated initial image $\boldsymbol{\hat{x}}_{0}$. Specifically, the forward diffusion process can be modeled as the solution to an Itô SDE:
@@
The ODE can compute the exact likelihood of any input data by leveraging the connection to neural ODEs Chen et al. (2018). To approximate the log density of noised data distribution $\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$ at each sampling step, a network $\boldsymbol{s}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t)$ is trained to learn the corresponding log density
@@ 2.2 DDIM Solver And Consistency Models
$\boldsymbol{\hat{x}}_{0}=\frac{\boldsymbol{x}_{t}-\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t)}{\sqrt{\alpha_{t}}}.$ (7)
@@ 2.3 Posterior Sampling in Diffusion Models
$\boldsymbol{x}_{t-1}=f\left(\boldsymbol{x}_{t},\boldsymbol{\hat{x}}_{0},\boldsymbol{\epsilon}\right)+\eta\nabla_{\boldsymbol{x}_{t}}\left\|y-\mathcal{A}(\boldsymbol{\hat{x}}_{0})\right\|^{2}_{2},\ \ \boldsymbol{\epsilon}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}\right),$ (12)
@@ 3 Method
We propose a novel sampling process integrated with a tailored optimization procedure that incorporates the measurements $\boldsymbol{y}$ and Langevin dynamics to enhance the quality of image reconstruction and editing. The adopted SDE/ODE solver is based on DDIM, as described in Eq. 6. Denote $\boldsymbol{z}\sim\mathcal{E}(\boldsymbol{x}_{0})$, $\boldsymbol{z}\in\mathbb{R}^{p}$ where $\mathcal{E}$ is an encoder and $\boldsymbol{x}_{0}$ is an initial image. Our method operates in latent space and leverages the theory of the posterior sampling to correct the bias from the initial features and introduce the target concepts. The core insight is using $\boldsymbol{y}$ in the form of Gaussian distribution, estimated $\boldsymbol{z}_{0}$ and Langevin dynamics as the optimization terms to correct the errors of the sampling process. The importance of reconstruction and the algorithm are introduced specifically in (Sec. 3.1). The implementation details of the editing process are illustrated in detail (Sec. 3.2). PostEdit takes around 1.5 seconds and 18 GB memory costs on a single NVIDIA A100 GPU.
@@ 3.1 Posterior Sampling for Image Reconstruction
The quality of reconstruction is a crucial indicator for evaluating the editing capabilities of a method. To preserve the features of the background (areas unaffected by the target prompt), Mokady et al. (2023) introduces a technique for fine-tuning the text embedding to mitigate errors caused by the null text term, as demonstrated in App. A.14. However, this approach is time-consuming, and there is a pressing need to enhance editing performance. To address this challenge, we propose a method that enables a fast and accurate reconstruction and editing process.
@@
Specifically, there are four steps in our method: (1) We add noise to $\boldsymbol{z}_{0}$ following the DDPM noise schedule until $\boldsymbol{z}_{T}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. Unlike the iterative inversion process used in Mokady et al. (2023) (DDIM inversion), where multiple network inferences are required, the added noise here is directly sampled from $\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. As a result, this process adds random noise directly to the clean image in a single step, significantly reducing the computational time. (2) Existing SDE/ODE solvers, such as DDIM and LCM, are employed to estimate $\boldsymbol{\hat{z}}_{0}$. (3) To ensure that $\boldsymbol{\hat{z}}_{0}$ aligns more consistently with the background and target prompt features of the original image, it is optimized using two $L_{2}$ norm terms related to the defined measurement $\boldsymbol{y}$ and $\boldsymbol{\hat{z}}_{0}$ respectively. Additionally, Langevin dynamics is employed to avoid convergence to local optima. (4) Finally, by progressively applying the above process to update the mean and variance of the Gaussian distribution in the Step (1) according to a predefined time schedule, we obtain $\boldsymbol{z}_{0}^{*}$ with consistent initial features and accurate target characteristics respectively as $T$ converges to 00. The presented algorithm corresponding to the above process is shown in detail in Fig. 2 for image reconstruction or editing and Alg. A.15 for image reconstruction. The measurement $\boldsymbol{y}$, defined in Eq. 17, can be simplified as a masked observation. For a deeper understanding, we recommend referring to Kawar et al. (2021).
@@
Our method requires a large text-to-image diffusion model as input and we select Stable Diffusion (SD) Rombach et al. (2022). Given that SD is trained on the dataset containing billions of images, the generated result has strong randomness relating to the same prompt. Moreover, reconstructing an image with a given text prompt starting from $\mathcal{N}\left(\boldsymbol{0},\boldsymbol{I}\right)$ usually gets poor results for SD due to the bias caused by the unconditional term in CFG. Therefore, if we directly apply posterior sampling strategy shown in Eq. 12 to acquire $\boldsymbol{z}_{0}$ from $\boldsymbol{z}_{T}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ by leveraging SD to inference noise, $\boldsymbol{\hat{z}}_{0}$ differs greatly from the ground truth $\boldsymbol{z}_{0}$. Conversely, this sampling process has good reconstruction performance when it leverages the diffusion model trained on small datasets, for example, FFHQ Karras et al. (2019) and ImageNet Deng et al. (2009) as shown in Chung et al. (2023); Zhang et al. (2024a). We experimentally discover that the gap between these two kinds of models is the inconsistent layouts of each estimated image $\boldsymbol{\hat{z}}_{0}$, while the features of the target prompts are successfully introduced into the generated $\boldsymbol{\hat{z}}_{0}$. Specifically, the layouts of $\boldsymbol{\hat{z}}_{0}$ generated by the scores inferred by the networks trained on FFHQ and SD for intermediate timesteps are shown in App. A.4. Therefore, due to the editing and reconstruction trade-off issue, it is much more challenging for high-quality image editing and reconstruction by leveraging large text-to-image models.
@@
to calculate the scores towards to $\boldsymbol{z}_{0}$ straightly inspired by Chung et al. (2023) and Zhang et al. (2024a). The measurement settings for image reconstruction are listed in the App. A.2
@@
$\boldsymbol{z}_{t-1}\sim\mathbb{E}_{\boldsymbol{z}_{0}^{w}}{\mathcal{N}}\left(\boldsymbol{z}_{0}^{w},\sigma_{t-1}^{2}\boldsymbol{I}\right),$ (15)
@@
satisfies the time marginal distribution conditioned on $p\left(\boldsymbol{z}_{t-1}|\boldsymbol{y}\right)$, where $\boldsymbol{z}_{0}^{w}$ is obtained from Eq. 13. (Proof is shown in Appendix A.16)
@@
Propsition 2 ensures that $\boldsymbol{z}_{t-1}$ sampled from the Gaussian distribution (with mean $\boldsymbol{z}_{0}^{w}$ and variance $\sigma_{t-1}^{2}$) still satisfies the constraint of the posterior sampling Eq. 10. Therefore, we can present the following scheme to optimize the estimated $\boldsymbol{\hat{z}}_{0}$ and run Langevin dynamics Welling & Teh (2011):
@@
$\boldsymbol{z}_{0}^{(k+1)}=(1-w)\cdot\boldsymbol{z}_{0}^{(k)}+w\cdot\boldsymbol{z}_{in}-h\cdot\nabla_{\boldsymbol{z}_{0}^{(k)}}\left(\frac{\|z_{0}^{(k)}-z_{0}\|^{2}}{2\sigma_{t}^{2}}+\frac{\|\mathcal{A}\left(\boldsymbol{z}_{0}^{(k)}\right)-\boldsymbol{y}\|^{2}}{2m^{2}}\right)+\sqrt{2h}\boldsymbol{\epsilon}.$ (16)
@@
Here, $\mathcal{A}\left(\cdot\right)$ is identical to $\boldsymbol{P}\left(\cdot\right)$ as defined in Eq. 17 and $h$ is the step size. $\sigma_{t}$ and $m$ are hyper-parameters detailed in Appendix A.2. Additionally, $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. Fig. 15 and Fig. 16 in Appendix A.10 demonstrate that PostEdit can achieve high-quality reconstruction outcomes without requiring any tuning process. Eq. 16 is reasonable since the two terms that multiplied by the step size $h$ have the same descent direction towards to the ground truth $\boldsymbol{z}_{0}$. Additionally, Langevin dynamics is employed to search for solutions that achieve a global optimum. Since the Step (1) shown in Fig. 2 is different from the process of DDIM inversion from the initial image to noise, which involves adding noise $\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{z}_{t},t,c_{ini}\right)$ inferred by the network at each step Mokady et al. (2023) (where $c_{ini}$ represents the prompt describing the content of the initial image), PostEdit is much more efficient as mentioned before. Considering that no information related to the initial image is incorporated into the noised distribution, the term involving the measurement $\boldsymbol{y}$, as defined in Eq. 16, is introduced to correct errors in the initial features caused by the unconditional term in CFG. This ensures background consistency. All parameter settings are detailed in Appendix A.2. The detailed process of image reconstruction is outlined in Alg. 2 of Appendix A.15.
Algorithm 1 Posterior Sampling for Image Editing
1:  Require: Diffusion model $\boldsymbol{\epsilon}_{\theta}$, step size $h$, posterior sampling steps $N$, diffusion solver steps $n$, image $\boldsymbol{x}_{0}$, measurement $\boldsymbol{y}$, weight $w$, target prompt $c_{tgt}$ , coefficients of diffusion sampler $c_{skip}$ and $c_{out}$, encoder ${\mathcal{E}}$, decoder $\mathcal{D}$, noise schedule $\alpha(t)$, $\sigma(t)$, posterior sampler sequence $\{\tau_{i}\}_{i=0}^{N-1}$ and diffusion sampler sequence $\{t_{j}\}_{j=0}^{n-1}$.
2:  $\boldsymbol{z}_{0}\sim\mathcal{E}(\boldsymbol{x}_{0})$, $\boldsymbol{z}_{in}=\boldsymbol{z}_{0}$
3:  for $i=N-1$ to 00 do
4:     for $j=n-1$ to 00 do
5:        Sample $\boldsymbol{z}_{j}\sim\mathcal{N}\left(\alpha\left(t_{j}\right)\boldsymbol{z}_{0},\sigma^{2}\left(t_{j}\right)\boldsymbol{I}\right)$
6:        $\boldsymbol{z}_{0}=c_{\text{skip }}(t)\boldsymbol{z}_{j}+c_{\text{out }}(t)\left(\frac{\boldsymbol{z}_{j}-\sigma_{t}\epsilon_{\boldsymbol{\theta}}(\boldsymbol{z}_{j},\boldsymbol{c_{tgt}},t)}{\alpha_{t}}\right)$
7:     end for
8:     $\boldsymbol{z}_{0}^{0}=\boldsymbol{z}_{0}$
9:     for $k=0$ to $T-1$ do
10:        Sample $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$.
11:        $\boldsymbol{z}_{0}^{(k+1)}=(1-w)\cdot\boldsymbol{z}_{0}^{(k)}+w\cdot\boldsymbol{z}_{in}-h\cdot\nabla_{\boldsymbol{z}_{0}^{(k)}}\left(\frac{\|z_{0}^{(k)}-z_{0}\|^{2}}{2\sigma_{t}^{2}}+\frac{\|\mathcal{A}\left(\boldsymbol{z}_{0}^{(k)}\right)-\boldsymbol{y}\|^{2}}{2m^{2}}\right)+\sqrt{2h}\boldsymbol{\epsilon}$
12:     end for
13:     Sample $\boldsymbol{z}_{\tau_{i-1}}\sim\mathcal{N}(\boldsymbol{z}_{0}^{(T)},\sigma_{{\tau_{i-1}}}^{2}\boldsymbol{I})$.
14:     $\boldsymbol{z}_{0}=c_{\text{skip }}(t)\boldsymbol{z}_{\tau_{i-1}}+c_{\text{out}}(t)\left(\frac{\boldsymbol{z}_{\tau_{i-1}}-\sigma_{t}\epsilon_{\boldsymbol{\theta}}(\boldsymbol{z}_{\tau_{i-1}},\boldsymbol{c_{tgt}},t)}{\alpha_{t}}\right)$
15:  end for
16:  $\boldsymbol{x}_{0}=\mathcal{D}\left(\boldsymbol{z}_{0}\right)$
17:  Return $\boldsymbol{x}_{0}$